OpenStudy (firejay5):

What's the inverse function of y = -3?

geerky42 (geerky42):

x = -3

geerky42 (geerky42):

It is not function so...

OpenStudy (firejay5):

so does it have an inverse function

OpenStudy (mertsj):

You can write it as y = 0x-3 Then interchange the x and y x = 0y - 3 or x = -3 But x =-3 is a vertical line and it is not a function.

geerky42 (geerky42):

So no inverse function? This is what Firejay5 and I want to know...

OpenStudy (mertsj):

No. It does not have an inverse function. It has an inverse but the inverse is not a function.

OpenStudy (firejay5):

so how should I explain it? It's not an inverse, because

geerky42 (geerky42):

Well... The inverse of y = -3 is x = -3, but x = -3 is not function so no inverse function.

Directrix (directrix):

A constant function does not have an inverse. Read about it on page one at this link: http://tinyurl.com/cnb9xc4

OpenStudy (mertsj):

That is the first time I ever heard that. We know that the inverse is not a function but does that mean it doesn't have an inverse? The function and its inverse do have symmetry about the line y = x.

OpenStudy (firejay5):

I am putting it, "No inverse, because a constant can't be an inverse

OpenStudy (firejay5):

would that be the right answer to the question

OpenStudy (mertsj):

http://www.purplemath.com/modules/invrsfcn2.htm

OpenStudy (kropot72):

So y = -3 has an inverse, however that inverse is not a function.

OpenStudy (firejay5):

@kropot72 the question that it asked me to do for each problem. Find the inverse for each one and graph it, so I can't graph it right

OpenStudy (kropot72):

Well you could graph the inverse ( x = -3 ) and make a note that it is not the graph of a function.

OpenStudy (firejay5):

what about y = -3

OpenStudy (kropot72):

The question appears to just require you to graph the inverse function and not the original function ( y = -3 )

OpenStudy (firejay5):

I meant to add graph the function and its inverse

geerky42 (geerky42):

Don't you know how to graph y = -3 ?

OpenStudy (firejay5):

yes

geerky42 (geerky42):

and x = -3 ? Just note that x = -3 is not inverse function.

OpenStudy (firejay5):

I will

OpenStudy (firejay5):

what's g(x) = 2x + 3 over 6

geerky42 (geerky42):

\[g(x) = \dfrac{1}{3}x + \dfrac{1}{2}\]

OpenStudy (firejay5):

\[\frac{ 2x + 3 }{ 3 }\]

geerky42 (geerky42):

2x + 3 over 6 or 2x + 3 over 3?

OpenStudy (firejay5):

@satellite73

geerky42 (geerky42):

Yup, it is g(x)=1/3 x+1/2

OpenStudy (firejay5):

when I graphed it it was only 1 line

geerky42 (geerky42):

It is because 1/3 x+1/2 is linear function.

OpenStudy (firejay5):

what about 2x + 3 over 6

geerky42 (geerky42):

Same thing. It is linear function.